Calculate the mean and standard deviation for the following data:


Neck circumference

(in inches)

12 12.5 13 13.5 14 14.5 15 15.5 16

Frequency

5 20 30 43 60 56 37 16 3


1
Expert's answer
2021-02-03T04:02:29-0500

To find the mean and standard deviation we will calculate the following table,



Here "A=" assumed mean "=14"

Now the formula for mean is, "\bar x = A+ \frac{\Sigma fd}{\Sigma f}"

"=14+\frac{3.5}{270}=14.0129"

"\therefore" The required mean is "=14.0129"

Again the formula for standard deviation is , "\sigma =\sqrt{\frac {\Sigma fd^2}{\Sigma f}-(\frac {\Sigma fd}{\Sigma f})^2}"

"=\sqrt{\frac {204.75}{270}-(\frac {3.5}{270})^2}"

"=0.8707"

"\therefore" The required standard deviation is "=0.8707"